OpenStudy (neonclawzz):

PLEASE PLEASE PLEEASE HELP! WILL MEDAL AND FAN! Form the perfect square trinomial in the process of completing the square. What is the value of c? x^2 + 3x + c = 7/4 + c c =

OpenStudy (ineedhelplz):

To complete the square you take the b term of the equation ax^2 +bx+c, in this case 3, and divide it by 2 then square it. so, the c term = (3/2)^2 = 9/4

OpenStudy (neonclawzz):

Thank you so much!

OpenStudy (ineedhelplz):

You're welcome!